OpenStudy (anonymous):

Exponential Growth. The population of a city 10 years ago was 45,600. Since then the population has increased at a steady rate each year. If the population is currently 64,800, find the annual rate of growth for this city. Please show work! For this problem the answer is 3.6%.

OpenStudy (anonymous):

i think you are supposed to solve \[64800=45600e^{10k}\] for k

OpenStudy (anonymous):

let's see if it works: \[\frac{27}{19}=e^{10k}\] \[\ln(\frac{27}{19})=10k\] \[k=\frac{\ln(\frac{27}{19})}{10}=.035\]\

OpenStudy (anonymous):

close was this the method, or where you supposed to do something else?

OpenStudy (anonymous):

Since the rate is steady I think you should do Y=a(1+r)^t

OpenStudy (anonymous):

exponential growth rate is very steady, but we can try it this way too

OpenStudy (anonymous):

I figured it out

OpenStudy (anonymous):

\[64,800=54,600(1+r)^{10}\] and solve for \(r\) \[\frac{27}{19}=(1+r)^{10}\] \[1+r=\sqrt[10]{\frac{27}{19}}\] \[1+r=1.0358\] \[r=.0358\approx 3.6\%\]